OpenStudy (anonymous):

What is the fourth term in this sequence? tn=3.2n-5

OpenStudy (anonymous):

put n=4 and solve

OpenStudy (anonymous):

7.8 i guess

OpenStudy (anonymous):

tn=3.2n-5 so t4 = 3.2(4) - 5 = 12.8 - 5 = 7.8

OpenStudy (anonymous):

true